OpenStudy (anonymous):

time (hours): 2, 4, 6, 8 distance (miles): 102, 153, 204, 255 what is the rate of change?

OpenStudy (anonymous):

segment by segment 0 to 2 hours rate = (102-0)/2 = 51 2 to 4 (153-102)/2= 51/2 4 to 6 (204-153)/2 = 51/2 6 to 8 (255-204)/2 = 51/2 first interval is different from the others?